OpenStudy (anonymous):

(-6x^5yz^7)(9xy^3z^2) step by step please, thanks (:

OpenStudy (curry):

-54x^6y^4z^9

OpenStudy (anonymous):

-54x^6y^4z^9 Here is how... -6*9 = 54 x^5 * x = x^6 y * y^3 = y^2 z^7*z^2 = 2^9

OpenStudy (anonymous):

pk51 is right, in the explanation he or she made an error though.. y*y^3 = y^4

OpenStudy (anonymous):

I was just aboout to say lol

OpenStudy (katrinakaif):

Utk? Can you please come to chat? (by Curry's request) Thanks! =)